Doubts
Sign in
12thPass logo
New chatPYQ LibraryDoubtsRank report
Sign in to see Recents
Sign in to save progress →
Doubts
Sign in

Probability question

2003 · Shift 0 · Q88
Guest · try one question or challengeGuest · try one question or challenge
  1. PYQ Library
  2. /JEE Main
  3. /Mathematics
  4. /Probability
  5. /2003 · Shift 0 · Q88

Probability question

2003 · Shift 0 · Q88

JEE MainMathematicsProbabilityMCQ+4 / −1
Events A,B,CA, B, CA,B,C are mutually exclusive events such that P(A)=3x+13,P(B)=1−x4P\left( A \right) = {{3x + 1} \over 3},P\left( B \right) = {{1 - x} \over 4}P(A)=33x+1​,P(B)=41−x​ and P(C)=1−2x2P\left( C \right) = {{1 - 2x} \over 2}P(C)=21−2x​ The set of possible values of xxx are in the interval.
  1. A
    [0,1]\left[ {0,1} \right][0,1]
  2. B
    [13,12]\left[ {{1 \over 3},{1 \over 2}} \right][31​,21​]
  3. C
    [13,23]\left[ {{1 \over 3},{2 \over 3}} \right][31​,32​]
  4. D
    [1 3,133]\left[ {{1 \ 3},{13 \over 3}} \right][1 3,313​]
PreviousNext

More from Probability

  • A problem in mathematics is given to three students A,B,C and their respective probability of solving the problem is 21​,31​ and 41​. Probability that the problem is solved is :2002 · MCQ
  • A and B are events such that P(A∪B)=3/4, P(A∩B)=1/4,P(A)=2/3 then P(A∩B) is :2002 · MCQ
  • Three distinct numbers are selected randomly from the set {1,2,3,…,40}. If the probability, that the selected numbers are in an increasing G.P., is nm​,gcd(m,n)=1, then m+n is equal to ​…2025 · Numerical
  •  Given three indentical bags each containing 10 balls, whose colours are as follows :  Bag I  Bag II  Bag III ​ Red 345​ Blue 231​ Green 534​…2025 · MCQ
  • If the probability that the random variable X takes the value x is given by P(X=x)=k(x+1)3−x,x=0,1,2,3…, where k is a constant, then P(X≥3) is equal to2025 · MCQ
  • The probability, of forming a 12 persons committee from 4 engineers, 2 doctors and 10 professors containing at least 3 engineers and at least 1 doctor, is2025 · MCQ
  • A box contains 10 pens of which 3 are defective. A sample of 2 pens is drawn at random and let X denote the number of defective pens. Then the variance of X is2025 · MCQ
  • A card from a pack of 52 cards is lost. From the remaining 51 cards, n cards are drawn and are found to be spades. If the probability of the lost card to be a spade is 5011​, then n is equal to ​ .2025 · Numerical